15t^2-29t-14\ 3t-7


answer:5t+2
need a full break down of how my book got this answer i go as far as 15t\3t+5t and -14t/-7=+2 but what happened to the -29t in this equation

Answer by stanbon(75887)

15t^2-29t-14\ 3t-7
Factor to get:
(3t-7)(5t+2)/(3t-7)
= 5t+2
-------------
Proof:
(3t-7)(5t+2)
Use FOIL:
F: 3t*5t = 15t^2
O: 3t*2 = 6t
I: -7*5t = -35t
L: -7*2 = -14
------------------
Combine to get:
15t^2 -29t - 14
===================
So the -29t comes from combining the Outer and the Inner products.
====================
Cheers,
Stan H.